# CasTGAN: Cascaded Generative Adversarial Network for Realistic Tabular Data Synthesis

### v1

@@ Abstract
Generative adversarial networks (GANs) have drawn considerable attention in recent years for their proven capability in generating synthetic data which can be utilized for multiple purposes. While GANs have demonstrated tremendous successes in producing synthetic data samples that replicate the dynamics of the original datasets, the validity of the synthetic data and the underlying privacy concerns represent major challenges which are not sufficiently addressed. In this work, we design a cascaded tabular GAN framework (CasTGAN) for generating realistic tabular data with a specific focus on the validity of the output. In this context, validity refers to the the dependency between features that can be found in the real data, but is typically misrepresented by traditional generative models. Our key idea entails that employing a cascaded architecture in which a dedicated generator samples each feature, the synthetic output becomes more representative of the real data. Our experimental results demonstrate that our model well captures the constraints and the correlations between the features of the real data, especially the high dimensional datasets. Furthermore, we evaluate the risk of white-box privacy attacks on our model and subsequently show that applying some perturbations to the auxiliary learners in CasTGAN increases the overall robustness of our model against targeted attacks.
Keywords Generative adversarial networks, tabular data, output validity
1 Introduction
Facilitating information and knowledge sharing within and between organisations is increasingly sought after for attaining growth and development. From a healthcare and medical standpoint, information exchange subsequently contributes to better understanding of diseases and risk factors, more intuitive prognosis by practitioners and effective treatment planning based on previously obtained knowledge. In the financial sector, sharing information between stakeholders leads to improved prediction of corporate bankruptcy and quicker identification of suspicious transaction behaviour that can be potentially linked to organised financial crime. For both fields, sharing the data which contains sensitive patient and client information is subject to GDPR regulations to maintain the confidentiality and privacy of such information. Therefore, institutions are continuously seeking new data anonymisation and synthetic data generation techniques for exchanging domain knowledge without exposing sensitive information.
@@ 1 Introduction
In contrast, GANs have been significantly less explored for tabular data generation. A tabular dataset typically comprises a mixture of continuous variables and categorical features. Tabular data is common in the medical and the financial domains where fields such age, gender, profession, and income can be commonly found in databases containing numerous records. As opposed to purely numerical data, representing datasets with categorical variables can be particularly difficult in presence of highly-dimensional and strongly correlated features. Furthermore, quantifying the validity of a synthetic tabular dataset can be practically impossible without closely inspecting every generated data sample and deciding whether to accept or reject each examined data record. Notwithstanding, there currently exists no straightforward and unified criteria for evaluating the validity of the output generated by tabular GANs.
To this end, we introduce CasTGAN, which is a generative network framework characterised by multiple generators connected sequentially; each of which is designed to generate a single feature. Meanwhile, a single discriminator validates the output of all the generators while being trained on the output of the final generator in the cascade. In addition, each generator is chained to a corresponding auxiliary learner in order to obtain more insightful losses specific to the individually generated features. This is motivated by the fact that it has been shown that adding more auxiliary classifiers can enhance the quality of the synthetic output images [28]. Therefore, we posit that CasTGAN aims to capture the highly correlated and hierarchical relationship between features, such that the synthetic output produced by our model closely resembles the real data while minimising the inconsistencies in the generated data. This is particularly important for applications where data is widely shared between professionals, and the slightest irregularities in the data can lead to undesired outcomes.
@@
  • Synthetic data evaluation: A new metric for quantifying the realisticness of the synthetic data when lacking the domain knowledge for the provided data, and extensively evaluate our framework and existing works.
@@
The remainder of this paper is structured as follows. In section 2 we present an overview of GANs and the types of GAN privacy attacks, while we further examine the relevant studies in section 3. Section 4 presents a discussion of CasTGAN and a detailed description of our model’s structure. In section 5, we discuss the experimental setup used in this work and the evaluation criteria. We demonstrate our results and a thorough analysis in section 6 and conclude in section 7.
2 Background
2.1 Generative Adversarial Networks
A GAN is characterised by a generator $G$ and a discriminator $D$ playing an adversarial game, where each component attempts to maximise its own benefit [10]. The generator receives a noise input sampled from a random distribution $z\sim p_{z}$ and learns to generate an output in the distribution $x\sim p_{g}$ that matches the structure of the unseen real data $x\sim p_{data}$. Meanwhile, the discriminator has access to the samples produced by the generator and the real data and learns to distinguish between its real and fake inputs. While the output generated by $G$ improves during training as a result of the loss it obtains from the discriminator, the discriminator also becomes increasingly clever in recognising the data produced by the generator. Subsequently, GANs are particularly challenging to train since it must be guaranteed that both the generator and the discriminator maintain their competitiveness without outperforming each other early in the training phase. In the classic GANs, the generator and the discriminator attempt to maximise their objective by minimising the Jenson-Shannon Divergence (JSD), however, using JSD does not guarantee the convergence of losses, hence leading to training instability [10].
@@ 2.1 Generative Adversarial Networks
WGAN-GP is increasingly becoming more prevalent than the classic GANs in applications such as image generation and tabular data generation, as it contributes to more stable learning. In addition, WGAN-GP minimises the effect of mode collapse - that is when the generator learns to "trick" the discriminator by producing a limited number of modes which the discriminator incorrectly classifies as real samples, instead of utilising the entire data feature space.
2.2 Privacy Attacks
@@ 3 Related Works
Tabular data is broadly used in regression and classifications tasks, which facilitates a growing interest in tabular data synthesis for machine learning applications, especially in domains with limited training data. Bayesian networks can be used for generating synthetic records by approximating the conditional probablity distribution from the data [47]. While the Bayesian networks can in practice be additionally used for exploring causal relationships between the independent variables, estimating the distributions is often built on simplifying assumptions on the data [9]. Meanwhile, tree-based methods were first utilised for generating partial synthetic data in [31], and has further explored in [40] where adversarial random forest has demonstrated comparable performance to deep learning techniques in terms of synthetic data quality.
@@
Deep neural networks have been widely studied for synthetic data generation, thanks to their capabilities for handling and approximating the distributions of large datasets. Variational auto-encoders (VAEs) [20] estimate the probabilistic distribution of by finding a lower-dimensional latent representation of the data. The application of VAEs has been extended to image data generation [39], oversampling of anomaly event data [14] and tabular data synthesis [42]. An underlying limitation with variational autoencoders is the assumption of the a simple parametric form of the latent space, which leads to a difficulty in capturing complex data distributions [13]. Invertible neural networks have also been proposed for tabular data synthesis through variants based on neural ordinary differential equations [22], copula flows [16] and normalizing flows for private tabular data generation [23].
@@
Within deep learning based generative models, GANs are favourable due to their ability to generate complex synthetic data in an adversarial and unsupervised setting [29]. table-GAN [30] is one of the earliest generative models for producing tabular synthetic output based on adversarial training. Using convolutional neural networks, table-GAN demonstrates that GANs unsurprisingly outperform anonymisation techniques while highlighting the potential privacy risks arising from membership attacks. Meanwhile, in medGAN [6] an autoencoder based generative model is developed for generating high-dimensional medical patient records while shedding light on the privacy risks attributed to the generated data [34]. A long-short term memory (LSTM) architecture for the generator was adopted in [43] demonstrating the potential of recurrent neural networks in synthetic data generation. A GAN approach based on an autoencoder and Kullback-Liebler divergence to tackle mode-collapse in GANs and producing synthetic output comparable to state-of-the-art methods. In CTGAN [42], conditional training of a GAN is carried out by instructing the generator and the discriminator to sample based on randomly selecting feature category in every training iteration, where a highly realistic tabular output can be observed from their evaluation. In [8], the authors propose cWGAN, which is a GAN-based oversampling technique focusing on the generation of samples belonging to the minority class in financial credit datasets. Zhao et. al [48] builds up on existing tabular GAN models to improve the representation of skewed distribution of numerical features of the synthetic output. Finally, Strelcenia et al. comprehensively review the existing tabular GAN literature and evaluate the studied variants for a class imbalance augmentation task [35].
@@
While there is no shortage of novelties in synthetic tabular data generation literature, an underlying challenge remains the proposal of evaluation techniques and criteria for quantifying the reliability and the statistical properties of the synthetic data. A further limitation is the sufficient analysis of data with hierarchical and interdependent variables. Therefore, our focus in this work is proposing a new framework that alleviates the two aforementioned deficiencies in the synthetic tabular data domain.
4 Methodology
@@
4.1 Model Architecture
The proposed CasTGAN framework is characterised by a cascade of multiple generators connected sequentially and coupled with auxiliary learners as illustrated in Figure 1.
Figure 1: The model architecture of CasTGAN.
The blocks in Figure 1 denoted with $G_{i}$ and $AL_{i}$, for $i=1,\ldots,M$, are respectively individual generators and auxiliary learners, one per feature $m_{i}$. Each generator $G_{i}$ focuses on generating its target feature using a primary neural network, and is laid out sequentially such that it obtains its inputs from a given noise vector whose components are standard Gaussian and i.i.d., and from the outputs of the previous generator (the only exception being the first generator which only takes a vector of random noise as its input). From a visual perspective, a generic generator $G_{i}$ may then be depicted as in Figure 2.
Figure 2: A close in visualization of one of the various generators in the cascaded layout.
Notation wise, the logic shown in Figure 2 is such that generator $G_{i}$ takes as inputs two objects: the useful outputs coming from $G_{i-1}$ (i.e., the vector $\check{G}_{i}(\phi_{i})$) and the noise vector $z$ (note that the same vector is fed to all the generators as depicted in Figure 1). The generator $G_{i}$ then produces one output, i.e., $G_{i}(\phi_{i})$, that may though be logically split in three distinct components: $Z_{i}$, that will be considered redundant information and that will not be used by the next generator; $\widehat{X}_{i}$, that is the target feature of generator $i$; and $\check{G}_{i-1}(\phi_{i-1})$, that is simply the information from the past generator that will be forwarded to the next one. This means that the output of $G_{i}$ can be formally presented as
@@ 4.1 Model Architecture
4.2 Auxiliary Learners Architecture
@@
On the other hand, the use of auxiliary learning for predicting the target variable given the data represents an alternative approach for capturing the characteristics of the data attributed to given target feature. It has been demonstrated that the auxiliary loss further stabilises the training process in comparison to conditional generation, and leads to a representation that it is independent of target label [38]. We observe that while auxiliary learners are traditionally embedded within the discriminator [28], we instead propose designing auxiliary learners as independent structures.
@@
In the CasTGAN, we craft M auxiliary learners $\{AL_{1},\ldots,AL_{M}\}$ for learning to predict the individual features. Due to its scalability on large datasets and the relatively fast convergence speed, we focus on building the auxiliary learners using the Light Gradient Boosting decision trees (LightGBM) [17], which we pre-train prior to the GAN training. An auxiliary learner $AL_{i}$ corresponding to feature $m_{i}$ is trained on $X_{\not\in i}$ in order to predict $X_{i}$. Following standard strategies for such tasks, for predicting the numerical features, the mean-squared error loss is used in the training of the auxiliary learners, whereas cross-entropy loss is used for predicting the categorical and binary variables. As with other decision tree based models, there is no need to one-hot encode the categorical features in $X_{\not\in i}$, but instead the categories are converted into integer encodings. Meanwhile the LightGBM auxiliary learners are capable of handling numerical features with extreme magnitudes, and therefore numerical features are not scaled for auxiliary training.
@@
As LightGBM models have low computational complexity and are generally fast to to train [17], assigning an auxiliary learner for every feature is a reasonable approach for representing the auxiliary loss $L_{AL}$ for predicting a feature given all the other features. It is worth noting that for the early auxiliary learners in the cascaded sequence $AL_{1}\text{ to }AL_{\lceil M/2\rceil}$, the generated data feature space $X_{\not\in i}$ is heavily dominated by redundant variables $Z$ which subsequently lead to increased auxiliary losses. However, these losses help the early generators in producing features that closely match the distributions of the training data. Meanwhile, the task for the later generators and auxiliary learners in the cascade becomes increasingly focused towards generating features that can be predicted from the initially generated target features $\check{G}_{i-1}\left(\phi_{i-1}\right)$.
@@
To ensure that the auxiliary losses $\{L_{AL_{1}},\ldots,L_{AL_{M}}\}$ do not overexceed the generator’s ones, there is the need for scaling down the losses from the auxiliary learners. In this paper we analyze the choice of performing this scaling down by means of constant coefficients $\lambda_{AL}$. In principle, $\lambda_{AL}$ could be a single scalar value that is applied to all the auxiliary losses. However, we consider a vector of auxiliary loss coefficients $\{\lambda_{AL_{1}},\ldots,\lambda_{AL_{M}}\}$ since it is fundamentally important for the early generators to generate variables that conform to the original feature distributions, since this will prompt the next generators in the cascade to effectively learn the feature correlations. Though this a hyperparameter, we set $\lambda_{AL_{1}}=0.75$ and $\lambda_{AL_{M}}=0.10$ while the auxiliary coefficients in between are linearly and equidistantly scaled in the $[0.75,0.10]$ range in our experiments. The same auxiliary setting is applied for all the datasets in this work.
4.3 Training Data Transformation

### v2

@@ Abstract
Generative adversarial networks (GANs) have drawn considerable attention in recent years for their proven capability in generating synthetic data which can be utilised for multiple purposes. While GANs have demonstrated tremendous successes in producing synthetic data samples that replicate the dynamics of the original datasets, the validity of the synthetic data and the underlying privacy concerns represent major challenges which are not sufficiently addressed. In this work, we design a cascaded tabular GAN framework (CasTGAN) for generating realistic tabular data with a specific focus on the validity of the output. In this context, validity refers to the the dependency between features that can be found in the real data, but is typically misrepresented by traditional generative models. Our key idea entails that employing a cascaded architecture in which a dedicated generator samples each feature, the synthetic output becomes more representative of the real data. Our experimental results demonstrate that our model is capable of generating synthetic tabular data that can be used for fitting machine learning models, as CasTGAN’s classification performance only falls under the real training data’s PR-AUC score by 4.88% on average for classification datasets, and exhibits an average reduction of the real training data’s $R^{2}$ score by 0.139 for regression datasets. In addition, our model captures well the constraints and the correlations between the features of the real data, especially the high dimensional datasets. Assessing the generation of invalid records demonstrates that CasTGAN reduces the number of invalid data observations by up to 622% in comparison to the second best performing baseline tabular GAN model. Furthermore, we evaluate the risk of white-box privacy attacks on our model and subsequently show that applying some perturbations to the auxiliary learners in CasTGAN increases the overall robustness of our model against targeted attacks.
Index Terms:
Generative adversarial networks, output validity, privacy attacks, tabular data.
[…]
=-21pt
I Introduction
Facilitating information and knowledge sharing within and between organisations is increasingly sought after for attaining growth and development. From a healthcare and medical standpoint, information exchange subsequently contributes to better understanding of diseases and risk factors, more intuitive prognosis by practitioners and effective treatment planning based on previously obtained knowledge [1]. In the financial sector, sharing information between stakeholders leads to improved prediction of corporate bankruptcy and quicker identification of suspicious transaction behaviour that can be potentially linked to organised financial crime [2]. For both fields, sharing the data which contains sensitive patient and client information is subject to the European Union’s General Data Protection Regulation (GDPR) [3] to maintain the confidentiality and privacy of such information. Therefore, institutions are continuously seeking new data anonymisation and synthetic data generation techniques for exchanging domain knowledge without exposing sensitive information.
@@ I Introduction
In contrast, GANs have been significantly less explored for tabular data generation. A tabular dataset typically comprises a mixture of continuous variables and categorical features [9]. Tabular data is common in the medical and the financial domains where fields such age, gender, profession, and income can be commonly found in databases containing numerous records. As opposed to purely numerical data, representing datasets with categorical variables can be particularly difficult in presence of highly-dimensional and strongly correlated features. Furthermore, quantifying the validity of a synthetic tabular dataset can be practically impossible without closely inspecting every generated data sample and deciding whether to accept or reject each examined data record. Correspondingly, invalidity refers to the semantically incorrect representation of the features, where the interdependencies between some data features are not correctly modelled. This constitutes a challenge when synthetic data is harnessed for knowledge exchange, as semantically incorrect data can lead to misinterpretation and flawed understanding of the data, hence disparaging the effect of facilitating knowledge sharing. Notwithstanding, there currently exists no straightforward and unified criteria for evaluating the validity of the output generated by tabular GANs [10].
To rectify the previously outlined limitations, we introduce CasTGAN, which is a generative network framework characterised by multiple generators connected sequentially; each of which is designed to generate a single feature. Meanwhile, a single discriminator validates the output of all the generators while being trained on the output of the final generator in the cascade. In addition, each generator is chained to a corresponding auxiliary learner in order to obtain more insightful losses specific to the individually generated features. This is motivated by the fact that it has been shown that adding more auxiliary classifiers can enhance the quality of the synthetic output images [11]. Therefore, we posit that CasTGAN aims to capture the highly correlated and hierarchical relationship between features, such that the synthetic output produced by our model closely resembles the real data while minimising the inconsistencies in the generated data. This is particularly important for applications where data is widely shared between professionals, and the slightest irregularities in the data can lead to undesired outcomes.
@@
  • Synthetic data evaluation: A new metric for quantifying the realistic-ness of the synthetic data when lacking the domain knowledge for the provided data, and extensively evaluate our framework and existing works.
@@
The remainder of this paper is structured as follows. In Section II, we present an overview of GANs and the types of GAN privacy attacks, while we further examine the relevant studies in Section III. Section IV presents a discussion of CasTGAN and a detailed description of our model’s structure. In Section V, we outline the experimental setup used in this work and the evaluation criteria. We demonstrate our results in Section VI and discuss our findings in Section VII. The paper is concluded in Section VIII.
II Background
II-A Generative Adversarial Networks
A GAN is characterised by a generator $G$ and a discriminator $D$ playing an adversarial game, where each component attempts to maximise its own benefit [4]. The generator receives a noise input sampled from a random distribution $z\sim p_{z}$ and learns to generate an output in the distribution $x\sim p_{g}$ that matches the structure of the unseen real data $x\sim p_{data}$. Meanwhile, the discriminator has access to the samples produced by the generator and the real data, and learns to distinguish between its real and fake inputs. While the output generated by $G$ improves during training as a result of the loss it obtains from the discriminator, the discriminator also becomes increasingly clever in recognising the data produced by the generator. Subsequently, GANs are particularly challenging to train since it must be guaranteed that both the generator and the discriminator maintain their competitiveness without outperforming each other early in the training phase. In the classic GANs, the generator and the discriminator attempt to maximise their objective by minimising the Jenson-Shannon Divergence (JSD), however, using JSD does not guarantee the convergence of losses, hence leading to training instability [4].
@@ II-A Generative Adversarial Networks
WGAN-GP is increasingly becoming more prevalent than the classic GANs in applications such as image generation and tabular data generation, as it contributes to more stable learning. In addition, WGAN-GP minimises the effect of mode collapse - that is when the generator learns to ”trick” the discriminator by producing a limited number of modes which the discriminator incorrectly classifies as real samples, instead of utilising the entire data feature space.
II-B Privacy Attacks
@@ III Related Works
Tabular data is broadly used in regression and classification tasks, which facilitates a growing interest in tabular data synthesis for machine learning applications, especially in domains with limited training data. Bayesian networks can be used for generating synthetic records by approximating the conditional probability distribution from the data [21]. While the Bayesian networks can in practice be additionally used for exploring causal relationships between the independent variables, estimating the distributions is often built on simplifying assumptions on the data [22]. Meanwhile, tree-based methods were first utilised for generating partial synthetic data in [23], and has further explored in [24] where adversarial random forest has demonstrated comparable performance to deep learning techniques in terms of synthetic data quality. However, the privacy-utility trade-offs for synthetic data generation using tree-based density estimators is not sufficiently explored.
@@
Deep neural networks have been widely studied for synthetic data generation, thanks to their capabilities for handling and approximating the distributions of large datasets. Variational auto-encoders (VAEs) [25] estimate the probabilistic distribution of by finding a lower-dimensional latent representation of the data. The application of VAEs has been extended to image data generation [26], oversampling of anomaly event data [27] and tabular data synthesis [28]. An underlying limitation with variational autoencoders is the assumption of the a simple parametric form of the latent space, which leads to a difficulty in capturing complex data distributions [29]. Invertible neural networks have also been proposed for tabular data synthesis through variants based on neural ordinary differential equations [30], copula flows [31] and normalizing flows for private tabular data generation [32]. A drawback of the invertible neural network based synthesis is that extensive hyperparameter tuning is needed to achieve satisfactory classification and regression performance.
TABLE I: Properties of some state-of-the-art tabular GANs, used as baselines later Section VI along with our proposed model, CasTGAN. When counting the number of datasets, we only consider real datasets as opposed to simulated datasets with known distributions and priors. While some methods conduct privacy analysis by measuring distance to closest records, this is notably different than evaluating the threat posed by privacy attacks.
Method Data types No. of real datasets Machine learning utility evaluation Univariate similarity evaluation Correlation analysis Validity evaluation Privacy attacks evaluation
table-GAN [33] mixed 4
medGAN [34] continuous 3
CTGAN [28] mixed 6
cWGAN [35] mixed 7
CTAB-GAN [36] mixed 5
CasTGAN [ours] mixed 6
@@
Within deep learning based generative models, GANs are favourable due to their ability to generate complex synthetic data in an adversarial and unsupervised setting [37]. table-GAN [33] is one of the earliest generative models for producing tabular synthetic output based on adversarial training. Using convolutional neural networks, table-GAN demonstrates that GANs unsurprisingly outperform anonymisation techniques while highlighting the potential privacy risks arising from membership attacks. Meanwhile, in medGAN [34] an autoencoder based generative model is developed for generating high-dimensional medical patient records while shedding light on the privacy risks attributed to the generated data. A long-short term memory (LSTM) architecture for the generator was adopted in [38] demonstrating the potential of recurrent neural networks in synthetic data generation. A GAN approach based on an autoencoder and Kullback-Liebler divergence to tackle mode-collapse was proposed in [39], in which high-quality synthetic output was produced, with comparable performance to state-of-the-art methods. In CTGAN [28], conditional training of a GAN is carried out by instructing the generator and the discriminator to sample based on randomly selecting feature category in every training iteration, where a highly realistic tabular output can be observed from their evaluation. In [35], the authors propose cWGAN, which is a GAN-based oversampling technique focusing on the generation of samples belonging to the minority class in financial credit datasets. Zhao et. al [36] builds up on existing tabular GAN models by employing convolutional neural network (CNN) and conditional vectors to improve the representation of skewed distribution of numerical features of the synthetic output. Finally, Strelcenia et al. [40] comprehensively review the existing tabular GAN literature and highlight that an underlying limitation of the tabular GAN-based studies is the lack of standarised evaluation metrics.
@@
Acknowledging the aforementioned, it is manifested that there is no shortage of novelties in synthetic tabular data generation literature. Nevertheless, an underlying challenge remains the proposal of evaluation techniques and criteria for quantifying the reliability and the statistical properties of the synthetic data. A further limitation is the sufficient analysis of data with hierarchical and interdependent variables. Therefore, our focus in this work is proposing a new framework that alleviates the two preceding deficiencies in the synthetic tabular data domain. A general overview of the properties, strengths and limitations of our model and the existing studies that we adopt as baselines is outlined in Table I.
IV Methodology
@@
IV-A Model Architecture
Figure 1: The model architecture of CasTGAN. The cascade $\vec{G}$ is composed of generators $G_{1},G_{2},\ldots,G_{M}$ sequentially lined up. The auxiliary learners $AL_{1},AL_{2},\ldots,AL_{M}$ are fitted on the real data, and are utilised by their respective generators for querying the generation of the data features $m_{1},m_{2},\ldots,m_{M}$. The cascade of generators $\vec{G}$ takes noise vector $z$ as input, while the discriminator $D$ is trained to distinguish the real data from the synthetic data. As depicted, generators $G_{1},G_{2},\ldots,G_{M-1}$ receive three losses: loss directly from the discriminator, loss backpropagated from the previous generator and the loss from the auxiliary learner. Meanwhile, $G_{M}$ is passed the loss from its auxiliary learner and the loss from the discriminator.
The proposed CasTGAN framework is characterised by a cascade $\vec{G}$ of multiple generators, $\vec{G}=\{G_{1},G_{2},\ldots,G_{M}\}$, in which generators $G_{1},G_{2},\ldots,G_{M}$ are connected sequentially and coupled with auxiliary learners $AL_{1},AL_{2},\ldots,AL_{M}$. Generator $G_{i}$ and auxiliary learner $AL_{i}$ for $i=1,2,\ldots,M$ are devoted for feature $m_{i}$ in the dataset, and the real data is used for fitting the auxiliary learners and the discriminator $D$. An illustration of the CasTGAN architecture is depicted in Figure 1.
As can be visualised from Figure 1, each generator $G_{i}$ focuses on generating its target feature using a primary neural network. The cascade of generators are laid out sequentially such that generator $G_{i}$ obtains its inputs from a given noise vector $z$ whose components are standard Gaussian and i.i.d., and from the outputs of the previous generator – the only exception being the first generator which only takes a vector of random noise as its input.
Figure 2: A close in visualization of generator $G_{i}$ in our GAN architecture. The structure in the figure is applicable to all generators in the cascaded layout, except for $G_{1}$, which receives only noise vector $z$ as input.
Notation-wise, generator $G_{i}$ takes as input $\phi_{i}$ two objects: the useful outputs coming from $G_{i-1}$ (i.e., the vector $\check{G}_{i-1}(\phi_{i-1})$) and the noise vector $z$ (note that the same vector is fed to all the generators as depicted in Figure 1). The generator $G_{i}$ then produces one output, i.e., $G_{i}(\phi_{i})$, that may though be logically split in three distinct components: $Z_{i}$, that will be considered redundant information and that will not be used by the next generator; $\widehat{X}_{i}$, that is the target feature of generator $i$; and $\check{G}_{i-1}(\phi_{i-1})$, that is simply the information from the past generator that will be forwarded to the next one. The input and the output of generator $G_{i}$ are shown in Figure 2.
Formally, the output of generator $G_{i}$ can be presented as:
@@ IV-A Model Architecture
IV-B Auxiliary Learners Architecture
@@
On the other hand, the use of auxiliary learning for predicting the target variable given the data represents an alternative approach for capturing the characteristics of the data attributed to given target feature. It has been demonstrated that the auxiliary loss further stabilises the training process in comparison to conditional generation, and leads to a representation that is independent of target label [42]. We observe that while auxiliary learners are traditionally embedded within the discriminator [11], we instead propose designing auxiliary learners as independent structures.
@@
In the CasTGAN, we craft M auxiliary learners $AL_{1},\ldots,AL_{M}$ for learning to predict the individual features. Due to its scalability on large datasets and the relatively fast convergence speed, we focus on building the auxiliary learners using the Light Gradient Boosting Machine (LightGBM) [43], which we pre-train prior to the GAN training. An auxiliary learner $AL_{i}$ corresponding to feature $m_{i}$ is trained on $X_{\not\in i}$ in order to predict $X_{i}$. Following standard strategies for such tasks, for predicting the numerical features, the mean-squared error loss is used in the training of the auxiliary learners, whereas cross-entropy loss is used for predicting the categorical and binary variables. As with other decision tree based models, there is no need to one-hot encode the categorical features in $X_{\not\in i}$, but instead the categories are converted into integer encodings. Meanwhile the LightGBM auxiliary learners are capable of handling numerical features with extreme magnitudes, and therefore numerical features are not scaled for auxiliary training.
@@
As LightGBM models have low computational complexity and are generally fast to to train [43], assigning an auxiliary learner for every feature is a reasonable approach for representing the auxiliary loss $L_{AL}$ for predicting a feature given all the other features. It is worth noting that for the early auxiliary learners in the cascaded sequence $AL_{1}\text{ to }AL_{\lceil M/2\rceil}$, the generated data feature space $\widehat{X}_{\not\in i}$ is heavily dominated by redundant variables $Z$ which subsequently lead to increased auxiliary losses. However, these losses help the early generators in producing features that closely match the distributions of the training data. Meanwhile, the task for the later generators and auxiliary learners in the cascade becomes increasingly focused towards generating features that can be predicted from the initially generated target features $\check{G}_{i-1}\left(\phi_{i-1}\right)$.
Figure 3: A detailed depiction of auxiliary learner $AL_{i}$ in CasTGAN. The training of the LightGBM model on the real data occurs prior to the training of the GAN model. Meanwhile, the synthetic data from the generator $G_{i}$ is queried against the auxiliary learner $AL_{i}$ during the training iterations of the GAN to compute the auxiliary loss of the generator’s target feature.
@@
To ensure that the auxiliary losses $L_{AL_{1}},\ldots,L_{AL_{M}}$ do not overexceed the generator’s ones, there is the need for scaling down the losses from the auxiliary learners. In this paper we analyze the choice of performing this scaling down by means of constant coefficients $\lambda_{AL}$. In principle, $\lambda_{AL}$ could be a single scalar value that is applied to all the auxiliary losses. However, we consider a vector of auxiliary loss coefficients $\lambda_{AL_{1}},\ldots,\lambda_{AL_{M}}$ since it is fundamentally important for the early generators to generate variables that conform to the original feature distributions, since this will prompt the next generators in the cascade to effectively learn the feature correlations. Though this a hyperparameter, we set $\lambda_{AL_{1}}=0.75$ and $\lambda_{AL_{M}}=0.10$ while the auxiliary coefficients in between are linearly and equidistantly scaled in the $[0.75,0.10]$ range in our experiments. The same auxiliary setting is applied for all the datasets in this work. The overall structure of auxiliary learner $AL_{i}$ is illustrated in Figure 3.
IV-C Training Data Transformation